OpenStudy (anonymous):

Find the slope-intercept form of each equation. x+3y=6

OpenStudy (anonymous):

Slope-Intercept Form \[y = mx + b\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

yeah thats the standard form of slope intercept form. so you got to get x+3y=6 into y = mx + b (aka slope intercept form)

OpenStudy (anonymous):

so 3y=mx+b?

OpenStudy (anonymous):

x+3y=6 Lets go step by step So first: subtract the x from both sides x+3y=6 -x -x --- ---- What do you get?

OpenStudy (anonymous):

3y=6-x

OpenStudy (anonymous):

YES! :) 3y = 6 - x BTW it is the same as 3y = -x + 6 Second: divide both sides by 3 3y = -x + 6 --- --- --- 3 3 3 what do you get?

OpenStudy (anonymous):

y=2-x

OpenStudy (anonymous):

or y=-x +2

OpenStudy (anonymous):

3y = -x + 6 -- -- --- 3 3 3 \[y = \frac{ -1{x} }{ 3 } + 2\] \[y = \frac{ -x }{ 3 } + 2\]

OpenStudy (anonymous):

why do you divide the middle one by 3 too?

OpenStudy (anonymous):

you were almost right :) but you forgot to divide the -x by 3

OpenStudy (anonymous):

because whatever you do to one side, you have to do to the other side. so basically, you can add, subtract, multiply, divide, but if you do it to one side you have to do it to the other side; hence both sides :)

OpenStudy (anonymous):

idk about that i already did that to the 6 on that side so i should be good

OpenStudy (anonymous):

3y = -x + 6 you divide by 3 on the left side to get the "y" by itself. 3y = -x + 6 -- --- --- 3 3 3 so you have to do it to the other side also.

OpenStudy (anonymous):

ok what now

OpenStudy (anonymous):

@some_someone

OpenStudy (shamim):

i wanna try to solve ur problem

OpenStudy (shamim):

given\[x+3y=6\]

OpenStudy (shamim):

\[3y=6-x\]

OpenStudy (shamim):

\[y=2-\frac{ 1 }{ 3 }x\]

OpenStudy (shamim):

\[y=-\frac{ 1 }{ 3 }x+2\]

OpenStudy (shamim):

now u hv to compare this equation with the standard eqution\[y=mx+c\]

OpenStudy (shamim):

so\[m=-\frac{ 1 }{ 3 }\]

OpenStudy (shamim):

\[m= slope\]

OpenStudy (shamim):

did u get it

OpenStudy (anonymous):

yes thanks